Simplify. (x^n/x^n-2)^3. Thanks for your help.
Found 2 solutions by Fombitz, rothauserc:
Answer by Fombitz(32388)
Answer by rothauserc(4718)

( (x^n) / (x^(n-2)))^3 = ( (x^3n) / (x^(3n-6))) = x^3n * (x^-3n) * x^6 = x^6
multiply exponents, denominator placed with numerator by using negative exponent, add like exponents